OpenStudy (anonymous):

Graph the exponential function f(x)=-(2/3) . To graph the function, plot the points on the graph with x-values -2,-1,0,1,and 2, draw all asymptotes.

OpenStudy (whpalmer4):

You didn't copy the problem successfully. Your formula is corrupted.

OpenStudy (anonymous):

Sorry it should be: Graph the exponential function \[f(x)=-(\frac{ 2 }{ 3 })^{x} \] . To graph the function, plot the points on the graph with x-values -2,-1,0,1,and 2, draw all asymptotes.

OpenStudy (ikram002p):



OpenStudy (anonymous):

how do i plot this?

OpenStudy (whpalmer4):

Make yourself a table of x and y values: x y -2 -(2/3)^-2 -1 -(2/3)^-1 0 -(2/3)^0 1 -(2/3)^1 2 -(2/3)^2 (you'll have to evaluate the expressions I wrote in the y column) Plot the points on your graph. Draw a smooth curve through them.

OpenStudy (whpalmer4):

Don't be worried if your graph looks different than @ikram002p's graph: he graphed \[y = (\frac{2}{3})^x\]not\[y = -(\frac{2}{3})^x\]

OpenStudy (anonymous):

so for the points I have: -2, -2.25, -1,-1.5, 0,-1,1,-0.67, 2,-0.44. is that right?

OpenStudy (whpalmer4):

Yes, here they are in both symbolic and numeric form: \[\begin{array}{ccc} x & y & y\\\hline\\ -2 & -\frac{9}{4} & -2.25 \\ -1 & -\frac{3}{2} & -1.5 \\ 0 & -1 & -1. \\ 1 & -\frac{2}{3} & -0.666667 \\ 2 & -\frac{4}{9} & -0.444444 \\ \end{array}\]

OpenStudy (ikram002p):

oh , sry i dint note the -ve sighn, how ever im she not he :o

OpenStudy (whpalmer4):

I wonder if your icon wasn't loading for me earlier today when I wrote that. It's not the sort of image that most guys would choose!

OpenStudy (ikram002p):

ya lol ! it would be weird if im he with this icon :o :P

OpenStudy (whpalmer4):

Indeed. Well, again, my apologies!

OpenStudy (ikram002p):

np, its ok ^^

OpenStudy (ikram002p):

http://www.wolframalpha.com/input/?i=-%282%2F3%29^x